OpenStudy (anonymous):

how many grams are in 23.0 moles of carbon?

OpenStudy (anonymous):

use periodic table and see whats 'molar mass' of carbon

OpenStudy (anonymous):

n = m / M so m = n * M and M is like sara said molar mass of carbon and n is number of moles

OpenStudy (anonymous):

1 mole = 12gm so 23 moles = 23*12=276gm